# June 13, 2024

### v2

@@ 5 Conclusion
This work presents a novel score-based black-box attack method named DifAttack++, which aims to hierarchically disentangle the adversarial and visual features from an image’s latent feature and perform black-box attacks via the disentangled feature space and cross domain. The main idea is to optimize the adversarial feature while keeping the visual feature unchanged until a successful AE is obtained. Experimental results demonstrate the superior attack efficiency of DifAttack++ in both close-set and open-set scenarios, as well as against a real-world API, meanwhile maintaining a satisfactory visual quality of AEs.
5.1 Acknowledgements

### v3

@@ 5 Conclusion
This work presents a novel query-efficient score-based black-box adversarial attack method named DifAttack++, which aims to hierarchically disentangle the adversarial and visual features from an image’s latent feature and perform black-box attacks via the disentangled feature space in cross-domain. The main idea is to optimize the adversarial feature while keeping the visual feature unchanged until a successful AE is obtained. Experimental results demonstrate the superior attack efficiency of DifAttack++ in both close-set and open-set scenarios, meanwhile maintaining a satisfactory visual quality of AEs.